OpenStudy (jaylah1212):

Function 1: y = 4x + 5 Function 2: The line passing through the points (1, 6) and (3, 10). Which of these functions has the greater rate of change? A) Function 1, because the slope is 5 and the slope of function 2 is 4. B) Function 1, because the slope is 4 and the slope of function 2 is 2. C) Function 2, because the slope is 7 and the slope of function 1 is 5. D) Function 2, because the slope is 5 and the slope of function 1 is 4.

OpenStudy (jaylah1212):

@rainbow_rocks03 @escamer @rainbowswirl

OpenStudy (rainbow_rocks03):

Sorry I dont understand this and im trying to get help myself too. Good Luck though

OpenStudy (jaylah1212):

@Jaynator495 @KendrickLamar2014 @LoveIt @LetsLearn2000 @littleelfie @gottennis121 @Qwertty123 @Error1603 @blackfireskull @briannaontimeforreal @xXiiBlueUnionXx

OpenStudy (welshfella):

the rate of change is given by the slpe of the line for the line y = mx + b the slope is the value of m

OpenStudy (welshfella):

* slope

OpenStudy (jaylah1212):

so it A

OpenStudy (welshfella):

How can it be A? what is the slope of the function y = 4x + 5?

OpenStudy (jaylah1212):

B

OpenStudy (welshfella):

the slope of 4x + 5 is 4 To find the slope of function 2 its the difference in y values / corresponding difference in x values

OpenStudy (jaylah1212):

IM SO CONFUSE @welshfella

OpenStudy (welshfella):

if the 2 points are (x1,y1) and (x2,y2) the slope is (y2 - y1) / (x2 - x1) for function 2 x1 = 1 ,x2 = 3 , y1 = 6 and y2 = 10.

OpenStudy (welshfella):

so this slope is (10-6) / (3 - 1)

OpenStudy (jaylah1212):

Can you just tell me A B C or D

OpenStudy (welshfella):

No I want you to learn how to do these types of problems now i#m sure you can work out (10-6) / (3 - 1) - work out the parenthese first then divide

OpenStudy (jaylah1212):

:( OK @welshfella

OpenStudy (jaylah1212):

@purple_pink

OpenStudy (jaylah1212):

I NEED HELP PLEASE

OpenStudy (victoriasushchik):

http://prntscr.com/aba4y0 so the answer is....

OpenStudy (jaylah1212):

SO it B @Victoriasushchik

OpenStudy (welshfella):

you only have to work out (10-6) / (3-1) to find the slope of function 2.

OpenStudy (jaylah1212):

ok so the answer is C

OpenStudy (welshfella):

10-6 = 4 3 - 1 = 2 4/2 = ?

OpenStudy (jaylah1212):

D

OpenStudy (welshfella):

you are guessing

OpenStudy (jaylah1212):

I NOT

OpenStudy (jaylah1212):

:( I really need help i will give you a medeal

OpenStudy (jaylah1212):

:( I really need help i will give you a medal

OpenStudy (victoriasushchik):

@jaylah1212 it is B you got it right!

OpenStudy (anonymous):

Rate of change (m) is: m=y2−y1x2−x1 In A the slope is: m=9−131−2=−4−1=4 In B the slope is: m=6−101−3=−4−2=2 So the answer is B

OpenStudy (welshfella):

surely you can divide 4 by 2??

OpenStudy (victoriasushchik):

fella I think she already got her answer

OpenStudy (jaylah1212):

Describe the relationship between the number 2 x 104 and 4 x 106. A) 4 x 106 is twice 2 x 104 B) 4 x 106 is 20 times more than 2 x 104 C) 4 x 106 is 200 times more than 2 x 104 D) 4 x 106 is 2000 times more than 2 x 104

OpenStudy (welshfella):

yea i know - but it seems she only wants the answer - not to learn how to get it. We shouldn't give answers out anyway.

OpenStudy (anonymous):

4 x 106 is 200 times more than 2 x 104

OpenStudy (victoriasushchik):

i didn't look at my older posts

OpenStudy (jaylah1212):

i think it a

OpenStudy (anonymous):

I just said it @jaylah1212

OpenStudy (victoriasushchik):

THE ANSWER IS \[B\]

OpenStudy (anonymous):

Actually its C

OpenStudy (jaylah1212):

Judy has a sugar cone and wants to know how many cubic inches of ice cream it will hold if it is filled completely to the top of the cone and no more. The cone has a height of 4.5 inches and a radius of 1.5 inches. A) 7.1 cubic inches B) 10.6 cubic inches C) 14.1 cubic inches D) 31.8 cubic inches

OpenStudy (jaylah1212):

wow I'm confuse i said A

OpenStudy (anonymous):

B

OpenStudy (victoriasushchik):

\(\color {#0cbb36}{\text{JAY }}\) THE ANSWER IS \[B\] \(\color{#0cbb34}{\text{DO YOU UNDERSTAND IT NOW?}}\)

OpenStudy (anonymous):

Judy has a sugar cone and wants to know how many cubic inches of ice cream it will hold if it is filled completely to the top of the cone and no more. The cone has a height of 4.5 inches and a radius of 1.5 inches. V = (1/3)pi.r^2.h = (3.1416*2.25*4.5)/3 = 1.0472*2.25*4.5 = (2.0944 + 0.2618)*4.5 = 2.3562*4.5 = 9.4248 + 1.1781 = 10.6029 cubic inches

OpenStudy (anonymous):

It would be B.

OpenStudy (jaylah1212):

Judy has a sugar cone and wants to know how many cubic inches of ice cream it will hold if it is filled completely to the top of the cone and no more. The cone has a height of 4.5 inches and a radius of 1.5 inches. A) 7.1 cubic inches B) 10.6 cubic inches C) 14.1 cubic inches D) 31.8 cubic inches

OpenStudy (anonymous):

We just said it was B.

OpenStudy (anonymous):

I explained the whole thing above ^^